# VideoRoPE: What Makes for Good Video Rotary Position Embedding?

### v1

@@ 1 Introduction
The original RoPE implementation (Vanilla RoPE) (Su et al., 2024) is designed for sequential 1D data like text. However, recent Video Large Language Models (Video LLMs) (Li et al., 2023; Lin et al., 2023a; Chen et al., 2024a; Maaz et al., 2024b; Zhang et al., 2024d; Wang et al., 2024d; Chen et al., 2024b; Zhang et al., 2024b) process video, which has a more complex spatio and temporal structure. As shown in Tab. 1, although several RoPE-based approaches (Gao et al., 2024; Wang et al., 2024b) have been proposed to support video inputs, these variants exhibit limitations and do not fully satisfy the following key characteristics:
@@
(2) Frequency Allocation. Previous approaches such as M-RoPE used in QWen2-VL (Wang et al., 2024b) employ 3D structure, dividing the feature dimensions into distinct subsets for ($t$, $x$, $y$) encoding, respectively. How to determine the optimal allocation of these dimension subsets, and consequently their associated frequencies  [^1] In RoPE, frequencies are determined by $\beta^{-2n/d}$, where $\beta$ is a constant, $n$ is the dimension index, $d$ is the total number of dimensions. Thus, choosing which dimensions represent $t$, $x$, and $y$ directly determines the frequencies used for each. are not well studied. Some previous work allocates the lower dimensions corresponding to the high frequency to represent the $t$. However, the temporal dimension $t$ is significantly tortured by periodic oscillation, and distant positions may have the same embeddings.
@@
(2) We introduce the challenging V-NIAH-D task to expose the drawbacks of current position embedding designs regarding frequency allocation. Our findings reveal that existing Video LLMs are easily misled to frequency-based distractors.
@@ 2 Related Work
Extending RoPE to Multi-Modal Data. Extending RoPE to multi-modal or Video LLMs typically follows two approaches. One approach directly applies standard RoPE, flattening visual tokens and treating text and visual tokens as a single 1D sequence. Although variants (e.g., TAD-RoPE (Gao et al., 2024)) introduce enhancements in indexing and attention mechanisms, these 1D RoPE variants overlook the spatiotemporal structure of video and inherent inter-modal differences (Su, 2024a, b; Wang et al., 2024b). In contrast, several studies have explored incorporating structural information to formulate the 2D/3D RoPE. For example, some previous works (Agrawal et al., 2024; Wang et al., 2024b) integrate RoPE-2D into visual encoders to improve spatial representation, particularly for resolution scaling. Based on the RoPE-Tie (Su, 2024a), M-RoPE (Wang et al., 2024b) used in QWen2-VL further generalizes RoPE to three dimensions to model both temporal and spatial dynamics. While effective, M-RoPE exhibits limitations, such as struggles with distractors in our V-NIAH-D task. This work presents a comprehensive analysis of the important characteristics essential for extending RoPE to video and proposes VideoRoPE according to our analysis.
@@ 3 Analysis
While the vanilla RoPE operates on 1D sequences, it can also be applied to higher-dimensional input by flattening the input into a 1-D sequence. However, the flattening process discards crucial neighborhood information, increases the sequence length, and hinders the capture of long-range dependencies. Therefore, preserving the inherent 3D structure is essential when adapting RoPE for video data. Some recent RoPE-variants (e.g., M-RoPE in Qwen2-VL (Wang et al., 2024b)) incorporate the $3$D structure. The corresponding relative matrix $\bm{A}_{(t_{1},x_{1},y_{1})}$ is computed as:
@@
Fig.  3 reveals unusual attention patterns in M-RoPE, despite its ability to locate the needle image but fails to answer the multi-choice question. According to the attention of M-RoPE, the needle is located primarily through vertical positional information, rather than temporal features. Thus, the temporal dimension fails to capture long-range semantic dependencies, focusing instead on local relationships. Conversely, the spatial dimensions exhibit a tendency to capture long-range rather than local semantic information. Lastly, the horizontal and vertical dimensions display distinct characteristics, with the vertical dimension exhibiting phenomena reminiscent of attention sinks (Xiao et al., 2023). These observations suggest that the performance decline primarily results from the sub-optimal frequency allocation designs of M-RoPE.
@@
Figure 4: (a) M-RoPE (Wang et al., 2024b) models temporal dependencies using the first 16 rotary angles, which exhibit higher frequencies and more pronounced oscillations. (b) In contrast, VideoRoPE models temporal dependencies using the last 16 rotary angles, characterized by significantly wider, monotonic intervals. Our frequency allocation effectively mitigates the misleading influence of distractors in V-NIAH-D. For a more detailed analysis, please refer to Appendix F.
@@ 4 VideoRoPE
Based on some previous research and the above analysis, we claim that a good RoPE design for Video LLMs, especially for long videos, should satisfy four requirements. The first requirement has been solved by RoPE-Tie (Su, 2024a) and the subsequent M-RoPE (Wang et al., 2024b). To solve the last three requirements and mitigate the performance decline observed in V-NIAH-D, we propose our VideoRoPE, comprising the following three key components.
Low-frequency Temporal Allocation (LTA). As shown in Eq. (2), the vanilla RoPE (Su et al., 2024) uses all dimensions to model the 1D position information. And as indicated in Eq. (4), M-RoPE (Wang et al., 2024b) uses dimensions to model temporal, horizontal, and vertical dimensions sequentially. However, previous frequency allocation strategies are suboptimal because different RoPE dimensions capture dependencies at varying ranges. As shown in Fig. 3, an interesting observation is that the local attention branch (as reported in (Han et al., 2024)) corresponds to lower dimensions, while the global branch (or attention sink, as in (Xiao et al., 2023)) corresponds to higher dimensions. To sum up, lower dimensions (higher frequency, shorter monotonic intervals, larger $\theta_{n}$) tend to capture relative distances and local semantics (Men et al., 2024; Barbero et al., 2024), while higher dimensions (lower frequency, wider monotonic intervals, smaller $\theta_{n}$) capture longer-range dependencies (Barbero et al., 2024).
@@
Figure 6: The 3D visualization for different position embedding. (a) The vanilla 1D RoPE (Su et al., 2024) does not incorporate spatial modeling. (b) M-RoPE (Wang et al., 2024b), while have the 3D structure, introduces a discrepancy in index growth for visual tokens across frames, with some indices remaining constant. (c) In contrast, our VideoRoPE achieves the desired balance, maintaining the consistent index growth pattern of vanilla RoPE while simultaneously incorporating spatial modeling.
@@ 5.1 Experimental Setup
Implementation Details. Using the aforementioned video training data, we fine-tune different modes that use different positional encoding strategies, such as the Vanilla RoPE (Su et al., 2024), Time-Aware Dual RoPE (TAD-RoPE) (Gao et al., 2024), M-RoPE (Wang et al., 2024b), and our VideoRoPE. All models are initialized with the Vision Transformer from Qwen2-VL-7B and LLM (Vanilla RoPE) from Qwen2-7B (Yang et al., 2024a). Our fine-tuning incorporates our VideoRoPE to process the spatiotemporal nature of the video data effectively. We adopt Qwen2-VL’s fine-tuning settings, processing each video at 2 fps with a maximum of 128 frames and dynamically adjusting the image resolution to maintain a consistent token count. However, to prevent memory overflow, we use a context window of 8192 tokens.
@@
Evaluation Benchmarks. We evaluate our approach using six video benchmarks, including tasks related to long video understanding, long video retrieval, and video hallucination. For long video understanding, we use LongVideoBench (Wu et al., 2024a) (8 seconds to 1 hour), MLVU (Zhou et al., 2024) (3 minutes to 2 hours), and Video-MME (Fu et al., 2024) (11 seconds to 60 minutes). For long video retrieval, we use Vision Needle-in-a-Haystack (V-NIAH) (Zhang et al., 2024d) and our proposed extension, Vision Needle-in-a-Haystack with Distractors (V-NIAH-D), which introduces distractor frames to increase the task difficulty. For video hallucination, we use VideoHallucer (Wang et al., 2024e), which evaluates the model’s ability to correctly answer both basic and hallucinated questions about video content. Details of these benchmarks can be found in Appendix B.
@@ 5.2 Results on Long Video Understanding
As shown in Tab. 2, we compare our VideoRoPE with existing RoPE variants (vanilla RoPE (Su et al., 2024), TAD-RoPE (Gao et al., 2024), and M-RoPE (Wang et al., 2024b)) across three prominent video understanding benchmarks. Our VideoRoPE consistently outperforms all baseline methods across these benchmarks, demonstrating its robustness and adaptability. Specifically, VideoRoPE achieves improvements of up to 2.91, 4.46, and 1.66 points (64k context length) over the M-RoPE baseline on LongVideoBench, MLVU, and Video-MME, respectively. These results emphasize the superior ability of VideoRoPE to effectively capture long-range dependencies and maintain performance across various challenging video data tasks.
@@ 5.5 Ablation Studies
Ablation Studies on Module Design. We conduct a series of ablation experiments on the modules introduced in Section 4. Through these experiments, we quantitatively evaluate the performance improvements, specifically analyzing their impact on the LongVideoBench and MLVU benchmarks. The experimental results are presented in Tab. 5. The baseline setting, M-RoPE (Wang et al., 2024b), achieves scores of 54.35 on LongVideoBench and 61.10 on MLVU (both using a 64k context length). By progressively integrating the DL (Diagonal Layout), LTA (Low-frequency Temporal Allocation), and ATS (Adjustable Temporal Spacing) modules, our method shows a continuous improvement in performance, achieving enhanced scores of 57.26 on LongVideoBench and 65.56 on MLVU (both using a 64k context length). These results demonstrate the effectiveness of our approach in leveraging spatial-temporal positional information. To refine the discussion on the allocation of $x$ and $y$ within LTA, we quantitatively examine the performance impact of interleaved versus sequential allocations. Additionally, we explore the optimal scaling factor in ATS by adjusting its values. Please refer to Appendix A.1 for more ablation studies on the scaling factor $\delta$ and the $x$/$y$ interleaved allocation.
@@ A.1 Supplementary Ablation Experiments
Table 7: Ablation Study on $x$, $y$ Allocation. VideoRoPE (Sequential) represents the sequential allocation of $x$ and $y$, following the pattern $x,x,x,\dots,y,y,y,\dots$ (similar to M-RoPE (Wang et al., 2024b)). VideoRoPE (Interleaved) represents the interleaved allocation, following the pattern $x,y,x,y,\dots$ (similar to Agrawal et al. (2024)).
@@ Appendix C More Related Works
Related Work on Video LLMs (Video Large Language Models) Video Large Language Models (Video LLMs) build upon the success of image-based vision-language models (VLMs) (Liu et al., 2023a; Zhang et al., 2023b; Dong et al., 2024; Zhang et al., 2024c; Zang et al., 2025; Chen et al., 2023; Chen & Xing, 2024; Liu et al., 2024c, b; Huang et al., 2024; Liu et al., 2024d; Xing et al., 2024), which align vision and language representations (Radford et al., 2021; Zhang et al., 2024a; Sun et al., 2023) but primarily focus on static images. Extending these models to video requires handling temporal dependencies (Xu et al., 2021; Lei et al., 2021; Bertasius et al., 2021; Huang et al., 2023) and long-form video understanding (Wang et al., 2024d; Chen et al., 2024b; Zhang et al., 2024d). Early video LLMs, such as Wang et al. (2023) and Li et al. (2023), leverage various Video Foundation Models (ViFMs), such as InternVideo (Wang et al., 2022), to extract video attributes, enabling LLM-based question answering. However, their ability to process video content is constrained by the limitations of ViFMs, restricting their effectiveness to short videos. To address this, Luo et al. (2023) introduces a Temporal Modeling Module, allowing end-to-end training of LLMs on video data. Building on this approach, Maaz et al. (2024a) further enhances spatiotemporal modeling to improve video comprehension. Meanwhile, Zhang et al. (2023a) and Lin et al. (2023b) integrate multiple modalities, such as audio and images, to enrich video understanding. These advancements lay the groundwork for processing long videos with greater accuracy. To extend Video LLMs’ capabilities to longer content, Lin et al. (2023c) first generates clip-level captions and then employs an LLM to integrate them into a comprehensive video caption, effectively representing the entire video. Various studies, such as those by Li et al. (2024b), Jin et al. (2023), Xu et al. (2024), and Zhang et al. (2025), explore different pooling strategies to reduce the number of video tokens, enabling LLMs to process longer videos more effectively. As the field progresses, there is a growing emphasis on long-form video understanding, exploring techniques such as streaming-based processing (Qian et al., 2025; Li et al., 2025), memory-augmented models (Qian et al., 2024; Ding et al., 2024a), and hierarchical representations (Wang et al., 2024f) to efficiently model extended temporal structures for tasks like event-level comprehension (Liu et al., 2024a) and video summarization (Chai et al., 2024).
@@
Related Work on Video Haystack Retrieval Originating from the Needle-In-A-Haystack task in Natural Language Processing (Kamradt, 2023; LangChain, 2024), Video haystack tasks aim to locate specific needle, the target information, within vast haystack, collections of video or multi-modal content (Zhang et al., 2024d; Wang et al., 2024c). In the video domain, VNBench (Zhao et al., 2024) first introduced a video haystack framework with diverse types of needles, such as subtitles, images, and video clips, specifically designed for retrieval tasks within a three-minute timeframe. V-NIAH (Zhang et al., 2024d) further advanced the field by extending retrieval tasks to long durations of up to one hour, providing tools for comprehensive evaluation. In a broader multi-modal domain, MMNeedle (Wang et al., 2024a) modeled the retrieval task as locating the exact coordinates of a sub-image within a larger multi-image haystack, while MM-NIAH (Wang et al., 2024c) introduced a setting where both haystack and needle could include images or text, emphasizing interleaved retrieval capabilities.

### v2

@@ 1 Introduction
The original RoPE implementation (Vanilla RoPE) (Su et al., 2024) is designed for sequential 1D data like text. However, recent Video Large Language Models (Video LLMs) (Li et al., 2023; Lin et al., 2023a; Chen et al., 2024a; Maaz et al., 2024b; Zhang et al., 2024d; Wang et al., 2024c; Chen et al., 2024b; Zhang et al., 2024b) process video, which has a more complex spatio and temporal structure. As shown in Tab. 1, although several RoPE-based approaches (Gao et al., 2024; Wang et al., 2024a) have been proposed to support video inputs, these variants exhibit limitations and do not fully satisfy the following key characteristics:
@@
(2) Frequency Allocation. Previous approaches such as M-RoPE used in QWen2-VL (Wang et al., 2024a) employ 3D structure, dividing feature dimensions into distinct subsets for ($t$, $x$, $y$) encoding, respectively. How to determine the optimal allocation of these dimension subsets, and their associated frequencies  [^1] In RoPE, frequencies are determined by $\beta^{-2n/d}$, where $\beta$ is a constant, $n$ is the dimension index, $d$ is the total number of dimensions. Thus, choosing which dimensions represent $t$, $x$, and $y$ directly determines the frequencies used for each. are not well studied. Some previous work allocates the lower dimensions corresponding to the high frequency to represent the $t$. However, the temporal dimension $t$ is significantly tortured by periodic oscillation, and distant positions may have the same embeddings.
@@
(2) We introduce the challenging V-NIAH-D task to expose the drawbacks of current position embedding designs regarding frequency allocation. We reveal that existing Video LLMs are easily misled to frequency-based distractors.
@@ 2 Related Work
Extending RoPE to Multi-Modal Data. Extending RoPE to multi-modal or Video LLMs typically follows two approaches. One approach directly applies standard RoPE, flattening visual tokens and treating text and visual tokens as a single 1D sequence. Although variants (e.g., TAD-RoPE (Gao et al., 2024)) introduce enhancements in indexing and attention mechanisms, these 1D RoPE variants overlook the spatiotemporal structure of video and inherent inter-modal differences (Su, 2024a, b; Wang et al., 2024a). In contrast, several studies have explored incorporating structural information to formulate the 2D/3D RoPE. For example, some previous works (Agrawal et al., 2024; Wang et al., 2024a) integrate RoPE-2D into visual encoders to improve spatial representation, particularly for resolution scaling. Based on the RoPE-Tie (Su, 2024a), M-RoPE (Wang et al., 2024a) used in QWen2-VL further generalizes RoPE to three dimensions to model both temporal and spatial dynamics. While effective, M-RoPE exhibits limitations, such as struggles with distractors in our V-NIAH-D task. This work presents a comprehensive analysis of the important characteristics essential for extending RoPE to video and proposes VideoRoPE according to our analysis.
@@ 3 Analysis
While the vanilla RoPE operates on 1D sequences, it can also be applied to higher-dimensional input by flattening the input into a 1-D sequence. However, the flattening process discards crucial neighborhood information, increases the sequence length, and hinders the capture of long-range dependencies. Therefore, preserving the inherent 3D structure is essential when adapting RoPE for video data. Some recent RoPE-variants (e.g., M-RoPE in Qwen2-VL (Wang et al., 2024a)) incorporate the $3$D structure. The corresponding relative matrix $\bm{A}_{(t_{1},x_{1},y_{1})}$ is computed as:
@@
Fig. 3 reveals unusual M-RoPE’s attention patterns, despite locating the needle image, it fails to answer the multi-choice question. According to M-RoPE’s attention, the needle is located primarily through vertical positional information, rather than temporal features. Thus, the temporal dimension fails to capture long-range semantic dependencies, focusing on local relationships. Conversely, the spatial dimensions capture long-range rather than local semantic information. Lastly, the horizontal and vertical dimensions display distinct characteristics, with the vertical dimension exhibiting phenomena reminiscent of attention sinks (Xiao et al., 2023). These suggest the performance decline primarily results from sub-optimal frequency allocation designs of M-RoPE.
@@
Figure 4: (a) M-RoPE (Wang et al., 2024a) models temporal dependencies using the first 16 rotary angles, which exhibit higher frequencies and more pronounced oscillations. (b) In contrast, VideoRoPE models temporal dependencies using the last 16 rotary angles, characterized by significantly wider, monotonic intervals. Our frequency allocation effectively mitigates the misleading influence of distractors in V-NIAH-D. For a more detailed analysis, please refer to Appendix F.
@@ 4 VideoRoPE
Based on some previous research and the above analysis, we claim that a good RoPE design for Video LLMs, especially for long videos, should satisfy four requirements. The first requirement has been solved by RoPE-Tie (Su, 2024a) and the subsequent M-RoPE (Wang et al., 2024a). To solve the last three requirements and mitigate the performance decline observed in V-NIAH-D, we propose our VideoRoPE, comprising the following three key components.
Low-frequency Temporal Allocation (LTA). As shown in Eq. (2), the vanilla RoPE (Su et al., 2024) uses all dimensions to model the 1D position information. And as indicated in Eq. (4), M-RoPE (Wang et al., 2024a) uses dimensions to model temporal, horizontal, and vertical dimensions sequentially. However, previous frequency allocation strategies are suboptimal because different RoPE dimensions capture dependencies at varying ranges. As shown in Fig. 3, an interesting observation is that the local attention branch (as reported in (Han et al., 2024)) corresponds to lower dimensions, while the global branch (or attention sink, as in (Xiao et al., 2023)) corresponds to higher dimensions. To sum up, lower dimensions (higher frequency, shorter monotonic intervals, larger $\theta_{n}$) tend to capture relative distances and local semantics (Men et al., 2024; Barbero et al., 2024), while higher dimensions (lower frequency, wider monotonic intervals, smaller $\theta_{n}$) capture longer-range dependencies (Barbero et al., 2024).
@@
Figure 6: The 3D visualization for different position embedding. (a) The vanilla 1D RoPE (Su et al., 2024) does not incorporate spatial modeling. (b) M-RoPE (Wang et al., 2024a), while have the 3D structure, introduces a discrepancy in index growth for visual tokens across frames, with some indices remaining constant. (c) In contrast, our VideoRoPE achieves the desired balance, maintaining the consistent index growth pattern of vanilla RoPE while simultaneously incorporating spatial modeling.
@@ 5.1 Experimental Setup
Implementation Details. Using the aforementioned video training data, we fine-tune different modes that use different positional encoding strategies, such as the Vanilla RoPE (Su et al., 2024), Time-Aware Dual RoPE (TAD-RoPE) (Gao et al., 2024), M-RoPE (Wang et al., 2024a), and our VideoRoPE. All models are initialized with the Vision Transformer from Qwen2-VL-7B and LLM (Vanilla RoPE) from Qwen2-7B (Yang et al., 2024a). Our fine-tuning incorporates our VideoRoPE to process the spatiotemporal nature of the video data effectively. We adopt Qwen2-VL’s fine-tuning settings, processing each video at 2 fps with a maximum of 128 frames and dynamically adjusting the image resolution to maintain a consistent token count. However, to prevent memory overflow, we use a context window of 8192 tokens.
@@
Evaluation Benchmarks. We evaluate our approach using six video benchmarks, including tasks related to long video understanding, long video retrieval, and video hallucination. For long video understanding, we use LongVideoBench (Wu et al., 2024a) (8 seconds to 1 hour), MLVU (Zhou et al., 2024) (3 minutes to 2 hours), and Video-MME (Fu et al., 2024) (11 seconds to 60 minutes). For long video retrieval, we use Vision Needle-in-a-Haystack (V-NIAH) (Zhang et al., 2024d) and our proposed extension, Vision Needle-in-a-Haystack with Distractors (V-NIAH-D), which introduces distractor frames to increase the task difficulty. For video hallucination, we use VideoHallucer (Wang et al., 2024d), which evaluates the model’s ability to correctly answer both basic and hallucinated questions about video content. Details of these benchmarks can be found in Appendix B.
@@ 5.2 Results on Long Video Understanding
As shown in Tab. 2, we compare our VideoRoPE with existing RoPE variants (vanilla RoPE (Su et al., 2024), TAD-RoPE (Gao et al., 2024), and M-RoPE (Wang et al., 2024a)) across three prominent video understanding benchmarks. Our VideoRoPE consistently outperforms all baseline methods across these benchmarks, demonstrating its robustness and adaptability. Specifically, VideoRoPE achieves improvements of up to 2.91, 4.46, and 1.66 points (64k context length) over the M-RoPE baseline on LongVideoBench, MLVU, and Video-MME, respectively. These results emphasize the superior ability of VideoRoPE to effectively capture long-range dependencies and maintain performance across various challenging video data tasks.
@@ 5.5 Ablation Studies
Ablation Studies on Module Design. We conduct a series of ablation experiments on the modules introduced in Section 4. Through these experiments, we quantitatively evaluate the performance improvements, specifically analyzing their impact on the LongVideoBench and MLVU benchmarks. The experimental results are presented in Tab. 5. The baseline setting, M-RoPE (Wang et al., 2024a), achieves scores of 54.35 on LongVideoBench and 61.10 on MLVU (both using a 64k context length). By progressively integrating the DL (Diagonal Layout), LTA (Low-frequency Temporal Allocation), and ATS (Adjustable Temporal Spacing) modules, our method shows a continuous improvement in performance, achieving enhanced scores of 57.26 on LongVideoBench and 65.56 on MLVU (both using a 64k context length). These results demonstrate the effectiveness of our approach in leveraging spatial-temporal positional information. To refine the discussion on the allocation of $x$ and $y$ within LTA, we quantitatively examine the performance impact of interleaved versus sequential allocations. Additionally, we explore the optimal scaling factor in ATS by adjusting its values. Please refer to Appendix A.1 for more ablation studies on the scaling factor $\delta$ and the $x$/$y$ interleaved allocation.
@@ A.1 Supplementary Ablation Experiments
Table 7: Ablation Study on $x$, $y$ Allocation. VideoRoPE (Sequential) represents the sequential allocation of $x$ and $y$, following the pattern $x,x,x,\dots,y,y,y,\dots$ (similar to M-RoPE (Wang et al., 2024a)). VideoRoPE (Interleaved) represents the interleaved allocation, following the pattern $x,y,x,y,\dots$ (similar to Agrawal et al. (2024)).
@@ Appendix C More Related Works
Related Work on Video LLMs (Video Large Language Models) Video Large Language Models (Video LLMs) build upon the success of image-based vision-language models (VLMs) (Liu et al., 2023a; Zhang et al., 2023b; Dong et al., 2024; Zhang et al., 2024c; Zang et al., 2025; Chen et al., 2023; Chen & Xing, 2024; Liu et al., 2024c, b; Huang et al., 2024; Liu et al., 2024d; Xing et al., 2024; Zhao et al., 2024a; Ding et al., 2025), which align vision and language representations (Radford et al., 2021; Zhang et al., 2024a; Sun et al., 2023) but primarily focus on static images. Extending these models to video requires handling temporal dependencies (Xu et al., 2021; Lei et al., 2021; Bertasius et al., 2021; Huang et al., 2023) and long-form video understanding (Wang et al., 2024c; Chen et al., 2024b; Zhang et al., 2024d). Early video LLMs, such as Wang et al. (2023) and Li et al. (2023), leverage various Video Foundation Models (ViFMs), such as InternVideo (Wang et al., 2022), to extract video attributes, enabling LLM-based question answering. However, their ability to process video content is constrained by the limitations of ViFMs, restricting their effectiveness to short videos. To address this, Luo et al. (2023) introduces a Temporal Modeling Module, allowing end-to-end training of LLMs on video data. Building on this approach, Maaz et al. (2024a) further enhances spatiotemporal modeling to improve video comprehension. Meanwhile, Zhang et al. (2023a) and Lin et al. (2023b) integrate multiple modalities, such as audio and images, to enrich video understanding. These advancements lay the groundwork for processing long videos with greater accuracy. To extend Video LLMs’ capabilities to longer content, Lin et al. (2023c) first generates clip-level captions and then employs an LLM to integrate them into a comprehensive video caption, effectively representing the entire video. Various studies, such as those by Li et al. (2024b), Jin et al. (2023), Xu et al. (2024), and Zhang et al. (2025), explore different pooling strategies to reduce the number of video tokens, enabling LLMs to process longer videos more effectively. As the field progresses, there is a growing emphasis on long-form video understanding, exploring techniques such as streaming-based processing (Qian et al., 2025a; Li et al., 2025), memory-augmented models (Qian et al., 2025b; Ding et al., 2024a), and hierarchical representations (Wang et al., 2024e) to efficiently model extended temporal structures for tasks like event-level comprehension (Liu et al., 2024a) and video summarization (Chai et al., 2024).
@@
Related Work on Video Haystack Retrieval Originating from the Needle-In-A-Haystack task in Natural Language Processing (Kamradt, 2023; LangChain, 2024), Video haystack tasks aim to locate specific needle, the target information, within vast haystack, collections of video or multi-modal content (Zhang et al., 2024d; Wang et al., 2024b). In the video domain, VNBench (Zhao et al., 2024b) first introduced a video haystack framework with diverse types of needles, such as subtitles, images, and video clips, specifically designed for retrieval tasks within a three-minute timeframe. V-NIAH (Zhang et al., 2024d) further advanced the field by extending retrieval tasks to long durations of up to one hour, providing tools for comprehensive evaluation. In a broader multi-modal domain, MMNeedle (Wang et al., 2025) modeled the retrieval task as locating the exact coordinates of a sub-image within a larger multi-image haystack, while MM-NIAH (Wang et al., 2024b) introduced a setting where both haystack and needle could include images or text, emphasizing interleaved retrieval capabilities.